La presentazione è in caricamento. Aspetta per favore

La presentazione è in caricamento. Aspetta per favore

CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 1 ECA Barrel ECC.

Presentazioni simili


Presentazione sul tema: "CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 1 ECA Barrel ECC."— Transcript della presentazione:

1 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 1 ECA Barrel ECC

2 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 2 Barrel Barrel in truck position in pit Infrastructure installation finished Electronics installation delayed by availability of components LV PS: Delays from manufacturer (MDI) First two proto failed load test (failing SMD capacitor) FEB: QPLL problem solved but still OTx problems 6 month delay: half should be recovered by new production/testing schedule All barrel FEBs by August (EC FEBs by November) Goal: install all electronics before moving to the IP? Cool down in IP: Start of cool down on 22/11/05 Barrel cold (will take 6 weeks) by end of December 2005 LArg filling beginning of Jan ‘06 Cryo ready for cold tests by mid Jan Then cold forever… Milano involvement: HV system commissioning

3 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 3

4 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 4 Endcap Calorimeters Endcap C Endcap A

5 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 5 Endcap C: Cold Test Cool-downWarm-up Normal working temperature (9 weeks)(7 weeks)

6 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 6 Endcap C: cold test results Cold test successfully performed Electromagnetic (EMEC): 6/31872 (0.19‰) dead channels: 5 EMEC (3 strips, 2 back) + 1 PS 16 HV problems (1408 lines): 13 sectors can be biased only on half of the semi-gaps (can be corrected SW: no acceptance loss) 2 sectors shorted together along  1 sector problems on both sides (x=0.2x0.2) at =2.3 (0.2%): could be operated at ~500V (half nominal voltage) Hadronic Endcap (HEC): 2/3072 dead channels (0.65‰) Forward (FCAL): 11 shorts (0.19%) Cryostat warm now: Start decabling and leave cryogenic system to Endcap A cold test Milano involvement: HV tests Time Domain Reflectometer (TDR) measurements: characterization of calibration lines (March ’05)

7 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 7

8 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 8 Endcap A Cold cover closed Feb 9 FCAL inserted, cabled and tested Welding of cold cover, purging of cold vessel, leak test, covers closed Cryostat displaced to cold test station ECA ready for cool down Cabling started last week Start of cool down: June 13 Cool down: 9 weeks Start of cold test: August 15 Cold test period: 6 weeks Milano involvement: HV set-up and tests TDR measurements

9 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 9

10 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 10 HV system status Milano is responsible of the HV control SW: Stable version since a while used for all tests up to now (CBT, Barrel/EC cold tests,…) Next step: integration with the rest of the DCS system Integration in the ATLAS DCS hierarchical system FSM: main tool for the implementation of the full control hierarchy in ATLAS, handle the states and transitions of the different parts of the detector. FSM: main tool for the implementation of the full control hierarchy in ATLAS, handle the states and transitions of the different parts of the detector. Iseg SW: Firmware and OPC server Iseg SW: Firmware and OPC server Unstable so far (v2.x, v3.x) Unstable so far (v2.x, v3.x) New version (v4.0) under test now, should fix bugs and instabilities New version (v4.0) under test now, should fix bugs and instabilities Will test new SW during ECA cold test starting in June Will test new SW during ECA cold test starting in June

11 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 11 Operator IF Data Viewer Alarm IF DCS_IS DIP Magnet CERN LHC DSS DAQ IS DAQ MRS DAQ RC CIC Global Control Station (GCS) Tile PixelSCTTRTLArMDTTGCRPC(CSC) LCS HEC LV LCS Temp LCS HV OPC LCS FE Crates HV/ LV PC Purity PVSS CAN DDC Local Control Stations (LCS) Subdetector Control Stations (SCS) Iseg HV modules Condition DB

12 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 12 Iseg HV modules HW responsibility: CERN, KTH, Wüppertal Modules still unstable During barrel cold test (August 2004) about half of the modules failed in going to 2kV: Problem traced to a bad batch of 10 nF/3 kV filter capacitors capacitors Capacitors replaced on all modules During Endcap C cold test (February ‘05) still several failures at V>2300 Volts: Failure of the voltage regulator Task force in action to investigate the problem and interact with the company: Three Milano people involved Meeting at CERN last week (also with Iseg people)

13 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 13 Iseg HV modules Tests performed in Milano, CERN, Iseg We built 8 load boxes (32 channels each) to do a systematic qualifying test of all modules at CERN: 1.5 F and 150 k in series on each channel to mimic a typical calorimeter HV channel Long duration tests Boxes at CERN since last week Iseg examined defective modules: opto-coupler failure in the regulation circuitry opto-coupler replaced (3 kV → 7 kV) on 30 modules But origin of fault not fully understood New modules at CERN since May 9: Modules under test w/ and w/o load boxes at maximum voltage (2500 V) So far so good Unforeseen and very urgent activity

14 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 14 CBT: list of data collected Low energy & linearity study: e - +  (no inner detector) E=1->9 GeV eta=0.2, 0.25, 0.35, 0.45, 0.55, 0.65 magnetic field ON for some runs 1/2/3 Al plate(s) in front of LARG: eta=0.4 LARG HV scan: e GeV eta= , 1500, 1000, 500, 250, 125 & 0 Volts on one side of the drift gap e GeV eta= , 50, 100, 150, 200, 400 & 700 Volts on both sides of the drift gap Phi-scan of sector 4 with dead electrodes: e GeV, eta=0.7 Fine eta scan 180 GeV with scintillator between LAr & Tiles e + &  180 GeV Fine eta scan 250 GeV e - &  250 GeV Magnet scan at eta=0.45,  250 GeV 25ns data (basically no physics data) protons 350 GeV: data with physics code in DSP: Material study with all subdetectors (e -, e + ) E=1->9 GeV E=20, , 150, 180 GeV eta=0.35, 0.4, 0.45, 0.5, 0.55, material between pixel & SCT, SCT & TRT, B-field ON/OFF Special studies for TRT, pions in magnetic field + lead absorber  20, 50, 100 GeV – B-field varied Vary B-field µ 150 GeV eta=0.45 e + 20, 50, 100 GeV eta=0.45 Study with muon chamber in front of LARG: e GeV eta=0.1,0.2,0.3,0.4,0.5,0.7,0.6 Photon run with tracker in: e GeV (e + ~= 150 GeV,  ~= 30GeV) Photon run with only LARG & TRT: e GeV (photon energies varied using B-field MBPL-H & V E  between 10 and 130 GeV) Non pointing photons - only LARG: Incoming e GeV -> E γ =50 GeV nominal eta=0.3 - displacement -20cm, 20, 40 and 60 cm

15 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 15 CBT: analysis Slow start-up of the analysis Delay runs collected in August-September Several software problems in switching from stand- alone framework to Athena Several problems with DB Lack of manpower End of February DB is finally full: Delay runs Master Waveform Calculation of optimal filtering coefficients can start single DAC calibration waves master waveform

16 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 16 Signal reconstruction and CBT RTM OFC (Milano Method) ready for all EMB module (all samplings) in DB Need to select the right OFC set (0,24) from TDC TDC needs to be calibrated Phase information from TDC is corrected with time offsets (found via iteration on data): Global: single value, related to general setup FEB: 1 value/FEB (16) Gain dependence TDC corrections finally available for all run periods Results for hi-p T electrons Energy scale still to be fixed Other corrections to be implemented: dead material, out-of-cone, leakage σ E / =0.97% E=250 GeV Expect:~0.7% = 182 GeV  / =0.87% E=180 GeV

17 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 17 CBT analysis Pedestal, ramps, delay & timing analysis finished and available in DB OFC: sets available and being tested: NTUPLE & ESD production using OFC starting First look at uniformity =0.4% Other studies: linearity (material studies) photon & muon runs pions Soon ready for combining track & calo for particle ID studies ±0.4%

18 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 18 Endcap Services Design phase completed in Milano Official drawings under finalization Milano is responsible for the installation of the Endcap services (ECC and ECA). Procurement of material starting (BNL, CERN) To minimize costs some materials will be shipped to Milano where they will be preassembled and tested Will bring in Milano mock- up/structure now in B180 help in the preparation of the various pieces Complicate service design All pieces are different from each other

19 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 19 Endcap services 13 electronic crates 4 differenti topologie di servizi: FCAL, EMEC, HEC/EMEC- special, Crates-feet Operazioni: PS: montaggio, cavi, schermi installazione water manifolds cooling FEB e PS cavi LV, cavi di trigger installazione delle box delle ELMB Installazione optical-boxes installazione parziale fibre TTC cavi di trigger test water lines ELMBLV PS Optical box HEC PS

20 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 20

21 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 21 Barrel Services

22 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 22 Calibration Calibrate em objects On DC2 samples Geant 4 Calibration hits: energy lost in dead materials (upstream, downstream, absorbers) Study various basic quantities: intrinsic response of the accordion Lateral containment Longitudinal leakage Energy lost between PS and strips... Goal is to parametrize these quantities as a function of a minimal set of observables Get to an em calibration scheme

23 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 23 Higgs analysis Complete analysis on inclusive channel H →  LOon DC1 samples ATLAS note under submission Also performed analysis on VBF channel: Low luminosity so far Higgs signal significance increases combining the two analysis: 10% improvement Next steps: Started validation of DC2 samples NLO analysis

24 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 24 E T miss,  reconstruction Study of Z →  → lep+had Control sample for reconstruction of ’s and for calibration of E T miss during commissioning: (M Z ) (E T miss ) 33k evts (Rome data) Digitized with noise in calorimeter Analysed from AOD and CBNT ROOT ntuple E T miss reconstructed from all calo cells in ||<5 (H1-style weights) Also from topoclusters Muons contribution (Moore muons) Cryostat correction Analysis cuts: PT(lep), PT(t) > 15 GeV <2.7 Next: Tune cuts Reconstruct also lepton background

25 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 25 Milestones 30/06/2004 Software per la ricostruzione del segnale con il metodo di Milano pronto al Test Beam combinato, compresa l`analisi delle linee di calibrazione e dei delay runs. 31/07/2004 Acquisizione misure con il Time Domain Reflectometer sul Barrel freddo e analisi preliminare 31/12/2004 Studio dell`uniformità del modulo 0 EM Barrel esposto al Test Beam combinato. 31/08/05 Installazione SW e inizio tests HV in USA15 31/08/05 Inizio misure HV sul Barrel nell`IP Completamento misure HV e TDR di ECA (in B180) Analisi dei dati del test beam in via di finalizzazione 31/10/05 Inizio integrazione dei servizi su EC-C in UX15 → OK (March 05) → OK → NO → scrittura software iniziata → settembre → settembre cold test, ottobre warm-up → novembre

26 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 26 Richieste finanziarie

27 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 27 Commissioning and Installation High Voltage activity: ECC cold test in B180 (fatto, parte si è dovuta fare nel 2005) ECA cold test in B180 (June- September) Inizio attivita’ nel pit: test HV sul barrel nell’IP, ECC (e ECA?) nel pit Integrazione col DCS di Atlas Attivita’ extra: HV task force, interazioni con Iseg (CERN, Dresda), endurance test dei moduli con i load boxes al CERN TDR measurements: ECC cold test in B180 (fatto) ECA cold test in B180 (August/September) Installazione Servizi ECC: preparazione al CERN e installazione servizi Team di 3 tecnici e supervisione del team al CERN da ottobre a dicembre Ad oggi: 1.5 mu HV e TDR su ECC A luglio: 4 mu 2 mu HV: ECA, load box, Iseg.5 mu: servizi ECC, disegni, mock-up Ad agosto/settembre: 5.5 mu 1.5 mu: finalizzazione HV e TDR ECA A fine anno: 10 mu 2 mu HV: test HV barrel, ECC e ECA nel pit, integrazione DCS 2.5 mu: servizi ECC nel pit Richiesta: 5 mu aggiuntivi missioni estere

28 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 28 Missioni Interne L’assegnazione di 14 k€ copre la ordinaria amministrazione e non tiene conto dei seguenti eventi: Liquid Argon week a Gargnano (12-16/09) organizzato da Milano Atlas Physics Workshop a Roma (6-11/06) Richiesta: 16 k€ aggiuntivi missioni interne 8 k€ per circa 8 persone a Gargnano 8 k€ per 5 persone a Roma

29 Back-up slides

30 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 30

31 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 31 Parameter analysis: Comparison with direct LC measurements (1) Systematic (mean) ~ 1% RMS ~ 2% Medium gain  extr =  meas for clarity Agreement with LC measured at cold is very good Reminder: 1% variation on LC leads to ~ 0.1% on peak reconstruction…

32 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 32 Parameter analysis: Calibration board parameters Medium gain L CB  8.8 H L CB  11.5 H R CB  2 

33 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 33 TDR measurements Length spread in each family 3  =1.5 ns (30cm). black TDR profile (TB ft=1,short line) green Z0=50 (Hspice simulation) red Z=Z0  2 (Hspice simulation)

34 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 34 TDR measurements cold pin carrier terminazione MB delay cavi di calibrazione baseplane vacuum cable warm cable warm pin carrier Blue: ionization signal Red: calibration signal

35 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 35 Ricostruzione del segnale: Strategia di calibrazione (2) Segnale di ionizzazione: Segnale di calibrazione: Funzione di trasferimento della catena di readout: Segnale di ionizzazione triangolare: Segnale di calibrazione “exponenziale”:

36 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 36 Ricostruzione del segnale: Strategia di calibrazione (3) Correzione per il punto di iniezione: LC Per completare l’equalizzazione delle celle, il guadagno di readout calcolato con il segnale di calibrazione… … va applicato a un segnale di ionizzazione che è stato corretto! Il segnale di ionizzazione triangolare generato nel gap di LAr è “normalizzato” quando corrisponde a un segnale di calibrazione unitario iniettato a livello della MB… Correzione per le diverse forme di corrente iniettata: T d, f step,  cali Le correzioni al segnale di ionizzazione sono implementate nella procedura di ricostruzione del segnale campionato…

37 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 37 A  Ricostruzione del segnale: “Optimal Filtering” L’ampiezza del segnale e il delay sono calcolati come combinazioni lineari dei campionamenti {S 1,…..,S N } I coefficienti a i e b i vengono calcolati a partire da: Forma “normalizzata” del segnale Autocorrelazione del noise Ampiezza del segnale delay Segnale “normalizzato” noise Il segnale è campionato ogni  t = 25 ns:

38 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 38 Ricostruzione del segnale: il metodo “Milano” 1.Obiettivo della procedura è l’estrazione di tutti i parametri utili alla predizione del segnale di ionizzazione dall’impulso di calibrazione  cali fit esponenziale sulla coda del segnale f step minimizazione  2 coda step response LCminimizazione  2 coda cosine response LCanalisi in frequenza FFT segnale 2.Il guadagno elettronico di ogni canale viene direttamente misurato attraverso una procedura di fit lineare (Master Waveform) ! OFC [ A/ADC ] 3.Il segnale di ionizzazione “normalizzato” viene predetto direttamente con una convoluzione numerica, senza l’utilizzo diretto di dati da segnali di fisica. concordano allo 0.5%

39 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 39 Ricostruzione del segnale: 32 samp. test-beam delay runs t (ns) y(t) = m(t) x DAC + DAC 0

40 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 40 Ricostruzione del segnale: Parametri dell’impulso di calibrazione P15 Middle FT0

41 CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 41 Ricostruzione del segnale: Parametro  0 = 1/LC Misura diretta FFT del segnale di calibrazione: blackraw transform redexp. correction blueshaper correction


Scaricare ppt "CSN1, 17 maggio 2005 Calorimetria EM a LAr di ATLAS 1 ECA Barrel ECC."

Presentazioni simili


Annunci Google